# Insights from leptohadronic modelling of the brightest blazar flare

### v1

@@ Appendix C Spectral-energy distributions for four other periods
Appendix D How do various values of proton acceleration efficiency affect the expected neutrino yield?
@@
To investigate the impact of our choice of the acceleration parameter $\xi_{\mathrm{acc}}=10^{7}$ for protons, the same as for electrons (obtained from fitting the electromagnetic SED), we show how the IceCube muon neutrino daily yield from 3C 454.3 changes if we assume a different value of $\xi_{\mathrm{acc}}$ for protons in Fig. 17 (left axis, black circles). For each value of $\xi_{\mathrm{acc}}$, we determine the maximum proton energy by finding the intersection of the proton acceleration timescale $t^{\prime\mathrm{acc}}_{p}$ (Eq. 37) with the proton energy-loss timescale (dominated by the photopion losses, see the purple dotted curve in Fig. 7). Then we repeat the procedure for finding $\rho_{p/e}$ (Eq. 40) as was described in Sect. 3.3 until the Bayes factor (Eq. 39) drops to $10^{-2}$. The corresponding values of $\rho_{p/e}$ are shown in the same plot in Fig. 17 (right axis, blue squares). Since at a fixed proton energy density, SEDs of secondary photons increase with increasing proton maximum energy (i.e. decreasing $\xi_{\mathrm{acc}}$), the overshooting of the X-ray data happens at a lower normalization of the proton SED. Hence, $\rho_{p/e}$ decreases when $\xi_{\mathrm{acc}}$ decreases. In Figs. 18 and 19, we show photon and neutrino SEDs for $\xi_{\mathrm{acc}}=5\times 10^{4}$ and $\xi_{\mathrm{acc}}=10$ respectively. One can see that the peak energy of the neutrino SED shifts to higher values with decreasing $\xi_{\mathrm{acc}}$ (compare also to Fig. 11). The value of $\xi_{\mathrm{acc}}=5\times 10^{4}$ is an optimal value in terms of maximizing neutrino yield. For $\xi_{\mathrm{acc}}\leq 10^{3}$, $\gamma\gamma$-initiated cascades start to dominate over $BH$-initiated cascades and $\rho_{p/e}$ must be significantly decreased not to overshoot the X-ray data while the energy of the neutrino SED continues shifting to higher values resulting in the lower expected IceCube neutrino yield. Values $\xi_{\mathrm{acc}}\leq 10^{3}$ are required to produce neutrinos with energies similar to the energy of the recently discovered $>$100 PeV neutrino KM3-230213A (The KM3NeT Collaboration, 2025).
@@
Figure 17: IceCube daily muon neutrino yield from 3C 454.3 averaged over MJD 55517–55524 for different values of the proton acceleration parameter $\xi_{\mathrm{acc}}$ (black circles, left ordinate axis) along with the UL on the proton-to-electron energy density ratio $\rho_{p/e}$ as a function of $\xi_{\mathrm{acc}}$ (blue squares, right ordinate axis). The error bars for neutrino yields represent the numerical uncertainty $\sim 40$% of normalization of spectra of secondaries in hadronic processes (Cerruti et al., 2024).
@@
Figure 18: The photon and neutrino SED with model curves for the “standing-feature” leptohadronic model with a fit to the observed SED averaged over MJD 55517–5524 for proton acceleration parameter $\xi_{\mathrm{acc}}=5\times 10^{4}$. Shaded regions correspond to the spread of proton luminosities $0\leq u^{\prime}_{p}\leq u^{\prime\mathrm{UL}}_{p}$, where $u^{\prime\mathrm{UL}}_{p}$ is the UL corresponding to the Bayes factor with respect to the best-fit leptonic model $BF=10^{-2}$ ($\approx 3\sigma$ Gaussian equivalent). The black solid curve with a downward arrow shows the IceCube one-year sensitivity (see Eq. 42) for the UL of 2.44 muon neutrinos between $E^{o\,\mathrm{min}}_{\nu}=10^{14}$ eV and $E^{o\,\mathrm{max}}_{\nu}=10^{17}$ eV.
@@
Appendix E On the discrepancy of the location of the $\gamma$-ray emitting region with some works
In contrast to the results of Bonnoli et al. (2011); Hu et al. (2015), who modelled SEDs averaged over the days around the flare MJD 55517–55524 (the closest date is MJD 55527) or the results of Sahakyan (2021) for the period we are studying in this work, we could not obtain a fit with the dissipation radius $x<R_{\mathrm{BLR}}$ lying within the BLR due to the $\gamma\gamma$ absorption as explained in Sect. 3.1.5. The different conclusions about the value of $x$ might be explained by different assumptions about the BLR size and/or the disc luminosity (see Table 2) resulting in non-zero but less-than-unity optical depth for the pair-production process at the threshold. Zhou et al. (2024) pointed out that the uncertainty of the BLR size can be of the order of half a magnitude, Abolmasov & Poutanen (2017) demonstrated that the uncertainty in the geometry of the BLR can result in significant uncertainty in the density and angular distribution of the BLR photon field in the blob frame, and Hopkins (2024) even argued that the Shakura-Sunayev disc (Shakura & Sunyaev, 1976) might be a poor model for the accretion flows around the SMBH. These uncertainties may be the reason for different conclusions about the location $x$ of the emission region for the 3C 454.3 flare of MJD 55517–55524. Our assumption $L_{\mathrm{AD}}=3\times 10^{46}$ erg/s is close to the estimations obtained from the observations of the broad lines of the BLR by Sbarrato et al. (2012); Pian et al. (2005) in the standard assumption of the BLR covering factor $\xi_{\mathrm{BLR}}=0.10$ (although this value can vary, Naddaf et al., 2021; Naddaf & Czerny, 2024, e.g.,), and the assumption about the BLR radius follows the standard phenomenological relation $R_{\mathrm{BLR}}\propto\sqrt{L_{\mathrm{AD}}}$ (Ghisellini & Tavecchio, 2009; Kaspi et al., 2007). A systematic study of the influence of the uncertainties of the BLR and AD parameters on the estimation of $x$ is beyond the scope of the current work.

### v2

@@ Appendix D Linear relation between integral $\nu$ and $\gamma$ fluxes
Appendix E How do various values of proton acceleration efficiency affect the expected neutrino yield?
@@
To investigate the impact of our choice of the acceleration parameter $\xi_{\mathrm{acc}}=10^{7}$ for protons, the same as for electrons (obtained from fitting the electromagnetic SED), we show how the IceCube muon neutrino daily yield from 3C 454.3 changes if we assume a different value of $\xi_{\mathrm{acc}}$ for protons in Fig. 20 (left axis, black circles). For each value of $\xi_{\mathrm{acc}}$, we determine the maximum proton energy by finding the intersection of the proton acceleration timescale $t^{\prime\mathrm{acc}}_{p}$ (Eq. 36) with the proton energy-loss timescale (dominated by the photopion losses, see the purple dotted curve in Fig. 12). Then we repeat the procedure for finding $\rho_{p/e}$ (Eq. 39) as was described in Sect. 5 until the Bayes factor (Eq. 38) drops to $10^{-2}$. The corresponding values of $\rho_{p/e}$ are shown in the same plot in Fig. 20 (right axis, blue squares). Since at a fixed proton energy density, SEDs of secondary photons increase with increasing proton maximum energy (i.e. decreasing $\xi_{\mathrm{acc}}$), the overshooting of the X-ray data happens at a lower normalization of the proton SED. Hence, $\rho_{p/e}$ decreases when $\xi_{\mathrm{acc}}$ decreases. In Figs. 21 and 22, we show photon and neutrino SEDs for $\xi_{\mathrm{acc}}=5\times 10^{4}$ and $\xi_{\mathrm{acc}}=10$ respectively. One can see that the peak energy of the neutrino SED shifts to higher values with decreasing $\xi_{\mathrm{acc}}$ (compare also to Fig. 13). The values of $10^{4}\lesssim\xi_{\mathrm{acc}}\lesssim 10^{5}$ are optimal in terms of maximising the neutrino yield. For $\xi_{\mathrm{acc}}\leq 10^{3}$, $\gamma\gamma$-initiated cascades start to dominate over BH-initiated cascades, and $\rho_{p/e}$ must be significantly decreased not to overshoot the X-ray data while the peak energy of the neutrino SED continues shifting to higher values resulting in the lower expected IceCube neutrino yield. Values $\xi_{\mathrm{acc}}\leq 10^{3}$ are required to produce neutrinos with energies similar to the energy of the recently discovered $\gtrsim$100 PeV neutrino KM3-230213A (KM3NeT Collaboration, 2025).
@@
Figure 20: IceCube daily muon neutrino yield from 3C 454.3 averaged over MJD 55517–55524 for different values of the proton acceleration parameter $\xi_{\mathrm{acc}}$ (black circles, left ordinate axis) along with the UL on the proton-to-electron energy density ratio $\rho_{p/e}$ as a function of $\xi_{\mathrm{acc}}$ (blue squares, right ordinate axis). The error bars for neutrino yields represent the numerical uncertainty $\sim 40$% of normalisation of spectra of secondaries in hadronic processes (Cerruti et al., 2024).
@@
Figure 21: The photon and neutrino SED with model curves for the “standing-feature” leptohadronic model with a fit to the observed SED averaged over MJD 55517–5524 for proton acceleration parameter $\xi_{\mathrm{acc}}=5\times 10^{4}$. Shaded regions correspond to the spread of proton luminosities $0\leq u^{\prime}_{p}\leq u^{\prime\mathrm{UL}}_{p}$, where $u^{\prime\mathrm{UL}}_{p}$ is the UL corresponding to the Bayes factor with respect to the best-fit leptonic model $BF=10^{-2}$ ($\approx 3\sigma$ Gaussian equivalent). The crimson solid curve with a downward arrow shows the IceCube one-year sensitivity (see Eq. 41) for the UL of 2.44 muon neutrinos between $E^{o\,\mathrm{min}}_{\nu}=10^{14}$ eV and $E^{o\,\mathrm{max}}_{\nu}=10^{17}$ eV.
@@
Appendix F MCMC fits to the spectral-energy distributions during the flare
To estimate the uncertainty of the parameter values describing the seven-day averaged SED (MJD 55517–55524) around the peak of the flare as well as the SEDs of each individual day within the period, we first run the MIGRAD routine using iMINUIT to obtain the values of parameters delivering the minimum of $\chi^{2}$. Then, to explore the parameter space around the obtained minimum, for each of the seven days and the seven-day-averaged period, we initialise a Monte-Carlo Markov Chain (MCMC) (Foreman-Mackey et al., 2013) with $24$ walkers. The initial guess for each walker for each parameter $w_{i}$ is chosen as $\left[\hat{w}_{i}\times(1\pm rN\{0,1\})\right]$, where $\hat{w}_{i}$ is the best-fit value of the $i$th parameter from the MIGRAD optimisation, $r=0.25$ (0.1 for MJD 55523–55524 due to a failed first attempt of the fit), and $N\{0,1\}$ stands for the standard normal distribution. The priors for the parameters are chosen to be flat within the following segments (see the description of the model parameters in Sect. 3.1):
  1. 1.
    $0.1\hat{L}_{e}^{\prime};10\hat{L}_{e}^{\prime}]$;
  2. 2.
    $0.5\hat{s}_{e};2\hat{s}_{e}]$;
  3. 3.
    $0.02\hat{E}_{e}^{\prime{\mathrm{min}}};3\hat{E}_{e}^{\prime{\mathrm{min}}}]$;
  4. 4.
    $0.333\hat{E}_{e}^{\prime{\mathrm{max}}};30\hat{E}_{e}^{\prime{\mathrm{max}}}]$;
  5. 5.
    $0.333\hat{R}_{b}^{\prime};3\hat{R}_{b}^{\prime}]$;
  6. 6.
    $0.2\hat{B}^{\prime};5\hat{B}^{\prime}]$;
  7. 7.
    $0.1\hat{\Gamma};2\hat{\Gamma}$];
  8. 8.
    $0.5\hat{x};3\hat{x}$].
The flat priors might bias sampling towards higher parameter values for those parameters that span more than one order of magnitude. To investigate the effect of the priors, for the seven-day-averaged dataset, we run a new MCMC fit changing the priors for parameters spanning more than one order of magnitude from uniform to log-uniform. We obtain results very close to the ones with the uniform priors, with all posterior parameter credible intervals remaining close to those with uniform priors, within $\pm 10$% variation. Thus, we conclude that the posterior distributions around the minimum obtained with iMINUIT are driven by the likelihood function and do not distinguish between uniform and log-uniform priors for the chosen number of MCMC steps. The results below are reported for the flat priors.
The best-fit values of the model parameters are defined as the medians of the marginal posterior one-dimensional parameter distributions obtained after running the MCMCs, while their uncertainties are defined as the differences between the median and the $16$th and $84$th percentiles252525following the tutorial at https://emcee.readthedocs.io/en/stable/tutorials/line. The best-fit parameter values with their uncertainties are presented in Fig. 8, while corresponding corner plots are shown in Figs. 23, 24, 25, 26, 27, 28, 29, 30. A qualitative description of the results is presented in Sect. 4.1.
Figure 23: The corner plot (Foreman-Mackey, 2016) obtained from the MCMC fitting (Foreman-Mackey et al., 2013) of the seven-day-averaged SED (MJD 55517–55524) with a “standing-feature” leptonic model. $t^{\prime}_{\mathrm{sim}}$ here is the same as for the one-day-averaged SEDs with $t^{o}_{\mathrm{av}}=1$ d (see Eq. 2). The corner plot is obtained from an MCMC containing $3,000$ steps per walker with $3\times 8=24$ walkers after combining chains from all the walkers (discarding $1,000$ initial steps of each walker as a burn-in) and applying thinning of $4$. Dashed vertical lines in the one-dimensional marginal histograms represent $16\%$, $50\%$, and $84\%$ percentiles of the posterior marginal distributions. Contour lines in the 2D marginal histograms represent $11.8\%$, $39.3\%$, $67.5\%$, and $86.4\%$ 2D containment regions.
Figure 24: The same as Fig. 23 but for MJD 55517–55518.
Figure 25: The same as Fig. 23 but for MJD 55518–55519.
Figure 26: The same as Fig. 23 but for MJD 55519–55520.
Figure 27: The same as Fig. 23 but for MJD 55520–55521.
Figure 28: The same as Fig. 23 but for MJD 55521–55522 and, instead, 2, 500 initial steps of each walker are discarded as a burn-in and no thinning is applied due to slower convergence of the fit.
Figure 29: The same as Fig. 23 but for MJD 55522–55523.
Figure 30: The same as Fig. 23 but for MJD 55523–55524 and, instead, $200$ initial steps of each walker are discarded as a burn-in and thinning of $5$ is applied due to a better value of $\chi^{2}_{\mathrm{red}}$ obtained in this case.